# Parallel transport on matrix manifolds and Exponential Action

### v1

@@ 3. Transposable spaces, trace and Frobenius forms, and a family of tractable metrics
$\begin{gathered}\frac{1}{\beta_{0}}\langle\mathrm{P}_{\mathtt{a}}\mathtt{b},\mathtt{c}\rangle_{\vec{\beta}}=\operatorname{Tr}\frac{1}{2}\left([\mathtt{b},\mathtt{a}]\mathtt{c}+(1+\beta)([\mathtt{a}_{\mathfrak{a}},\mathtt{b}]\mathtt{c}-[\mathtt{b}_{\mathfrak{a}},\mathtt{a}]\mathtt{c})\right)\\ -(1+\beta)\operatorname{Tr}\left(\frac{1}{2}\left([\mathtt{b},\mathtt{a}]\mathtt{c}_{\mathfrak{a}}+(1+\beta)([\mathtt{a}_{\mathfrak{a}},\mathtt{b}]\mathtt{c}_{\mathfrak{a}}-[\mathtt{b}_{\mathfrak{a}},\mathtt{a}]\mathtt{c}_{\mathfrak{a}})\right)\right)\\ =\operatorname{Tr}\frac{1}{2}\left(\mathtt{b}[\mathtt{a},\mathtt{c}]+(1+\beta)(\mathtt{b}[\mathtt{c},\mathtt{a}_{\mathfrak{a}}]-\mathtt{b}_{\mathfrak{a}}[\mathtt{a},\mathtt{c}])\right)\\ -(1+\beta)\operatorname{Tr}\left(\frac{1}{2}\left(\mathtt{b}[\mathtt{a},\mathtt{c}_{\mathfrak{a}}]+(1+\beta)(\mathtt{b}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]-\mathtt{b}_{\mathfrak{a}}[\mathtt{a},\mathtt{c}_{\mathfrak{a}}])\right)\right)\\ =\frac{1}{2}\operatorname{Tr}\mathtt{b}[\mathtt{a},\mathtt{c}]+\frac{1+\beta}{2}\operatorname{Tr}(\mathtt{b}[\mathtt{c},\mathtt{a}_{\mathfrak{a}}]-\mathtt{b}_{\mathfrak{a}}[\mathtt{a},\mathtt{c}]-\mathtt{b}[\mathtt{a},\mathtt{c}_{\mathfrak{a}}])\\ -\frac{(1+\beta)^{2}}{2}\operatorname{Tr}(\mathtt{b}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]-\mathtt{b}_{\mathfrak{a}}[\mathtt{a},\mathtt{c}_{\mathfrak{a}}]).\end{gathered}$
It remains to prove the following, as by eq. 10, all terms on the next to last line behave antisymmetrically when we change the positions of $\mathtt{b}$ and $\mathtt{c}$
$\operatorname{Tr}\mathtt{b}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]-\operatorname{Tr}\mathtt{b}_{\mathfrak{a}}[\mathtt{a},\mathtt{c}_{\mathfrak{a}}]+\operatorname{Tr}\mathtt{c}[\mathtt{b}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]-\operatorname{Tr}\mathtt{c}_{\mathfrak{a}}[\mathtt{a},\mathtt{b}_{\mathfrak{a}}]=0.$
The first and last term cancel, since $\operatorname{Tr}\mathtt{b}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]=\operatorname{Tr}\mathtt{b}_{\mathfrak{a}}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]$ as $\mathfrak{a}$ is a subalgebra, while $\operatorname{Tr}\mathtt{c}_{\mathfrak{a}}[\mathtt{a},\mathtt{b}_{\mathfrak{a}}]=\operatorname{Tr}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}]\mathtt{b}_{\mathfrak{a}}=\operatorname{Tr}[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}}]\mathtt{b}_{\mathfrak{a}}$ as $[\mathtt{c}_{\mathfrak{a}},\mathtt{a}_{\mathfrak{a}_{{}_{\perp}}}]\subset\mathfrak{a}_{{}_{\perp}}$. The middle two terms cancel similarly, switching $\mathtt{b}$ and $\mathtt{c}$. ∎
@@ 3.1. Riemannian metrics from the family
The case $\beta=-1$ is particularly simple, as we see the metric eq. 17 reduces to a multiple of the trace form $\beta_{0}\operatorname{Tr}\mathtt{h}\mathtt{g}$, the connection $\Gamma(\xi,\eta)$ reduces to the first two terms, the geodesic is $X\operatorname{expm}(tX^{-1}\xi)$, or just $\operatorname{expm}(t\xi)$ if $X=\operatorname{I}_{\mathcal{E}}$, the operator $\mathrm{P}_{\mathtt{a}}$ is just $-\frac{1}{2}\operatorname{ad}_{\mathtt{a}}$, and using eq. 8, the parallel transport is $X\operatorname{expm}(\frac{t}{2}\mathtt{a})X^{-1}\eta\operatorname{expm}(\frac{t}{2}\mathtt{a})$. For compact groups, this is known.
However, the trace form metric may not be Riemannian. The family of metrics in theorem 1 could be considered a deformation of the trace form using the Lie algebra $\mathfrak{a}$. For certain choices of the subgroup $\mathfrak{a}$ and parameters $\vec{\beta}$, the metric becomes Riemannian. The condition for this could be analyzed by studying the eigenvalues of $\mathcal{I}_{\vec{\beta}}$ above, which could be written as
(27) $\mathcal{I}_{\vec{\beta}}:\mathsf{T}\circ\left(\beta_{0}(\rm{I}_{\mathcal{E}}-\mathrm{p}_{\mathfrak{a}})-\beta_{1}\circ\mathrm{p}_{\mathfrak{a}}\right)=\left(\beta_{0}(\rm{I}_{\mathcal{E}}-\mathrm{p}_{\mathfrak{a}})-\beta_{1}\circ\mathrm{p}_{\mathfrak{a}}\right)\circ\mathsf{T}.$
Note that the transpose $\mathsf{T}$ has eigenvalues ±$\pm 1$ on the spaces of symmetric and antisymmetric matrices. Write $\mathcal{V}_{\mathrm{sym}}$ and $\mathcal{V}_{\mathrm{skew}}$ for the symmetric and antisymmetric components of a transposable subspace $\mathcal{V}\subset\mathcal{E}$, the possible eigenpairs of $\mathcal{I}_{\vec{\beta}}$ are $(\beta_{0},(\mathfrak{a}_{{}_{\perp}})_{\mathrm{sym}}),(-\beta_{0},(\mathfrak{a}_{{}_{\perp}})_{\mathrm{skew}}),(-\beta_{1},\mathfrak{a}_{\mathrm{sym}}),(\beta_{1},(\mathfrak{a})_{\mathrm{skew}})$. The Riemannian condition implies at least one of the first two, and at least one of the last two eigenspaces vanish, assuming the appropriate signs of $\beta_{0}$ and $\beta_{1}$. These are the possible cases
@@
    $(\mathfrak{a}_{{}_{\perp}})_{\mathrm{sym}}=\{0\},(\mathfrak{a})_{\mathrm{sym}}=\{0\}$. In this case $\mathfrak{g}_{\mathrm{sym}}=\{0\}$, or $\mathrm{G}$ is a compact Lie subgroup of $\mathrm{SO}(n)$. In this case $\beta_{0}<0,\beta_{1}>0$. An example is $\mathrm{SO}(n)$ itself in section 5.
@@
    $(\mathfrak{a}_{{}_{\perp}})_{\mathrm{sym}}=\{0\},(\mathfrak{a})_{\mathrm{skew}}=\{0\}$. For example, $\mathrm{G}$ could be a product of of a compact subgroup (for $\mathfrak{a}_{{}_{\perp}}$) and a diagonal group (for $\mathfrak{a}$). We need $\beta_{0}<0,\beta_{1}>0$.
@@
    $(\mathfrak{a}_{{}_{\perp}})_{\mathrm{skew}}=\{0\},(\mathfrak{a})_{\mathrm{sym}}=\{0\}$. This shows $\mathfrak{a}=\mathfrak{g}_{\mathrm{skew}}$. An example is $\operatorname{GL}(n)$ in [28, 18], reviewwed under our point of view with parallel transport in section 4. The case of $\mathrm{Sp}(n)$ in [6] could be treated similarly. In this case $\beta_{0}>0,\beta_{1}>0$. Except for the case $\mathfrak{g}_{\mathrm{sym}}=\{0\}$ already considered, the trace form is nondefinite, it is positive on $\mathfrak{g}_{\mathrm{sym}}$ and negative on $\mathfrak{g}_{\mathrm{skew}}$. The metric in eq. 17 modify the trace form on the antisymmetric part by a negative multiplier $-\beta_{1}$. This is also known in [3].

### v2

@@ 4.2. Deformation using an intertwining operator on a subalgebra
$\begin{gathered}\langle(\mathtt{X},\hat{\mathcal{B}}\mathcal{P}_{\mathfrak{a}}\mathtt{X}),(\mathtt{X},\hat{\mathcal{B}}\mathcal{P}_{\mathfrak{a}}\mathtt{X})\rangle_{\tilde{\mathcal{G}},\vec{\beta}}=\beta_{0}\langle\mathtt{X},\mathtt{X}\rangle-\langle\beta_{0}\mathcal{B}(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})^{-1}\hat{\mathcal{B}}\mathcal{P}_{\mathfrak{a}}\mathtt{X},\hat{\mathcal{B}}\mathcal{P}_{\mathfrak{a}}\mathtt{X}\rangle\\ =\beta_{0}\langle\mathtt{X},\mathtt{X}\rangle-\langle\beta_{0}\mathcal{B}(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})^{-1}(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{B}^{-1}\mathcal{P}_{\mathfrak{a}}\mathtt{X},(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{B}^{-1}\mathcal{P}_{\mathfrak{a}}\mathtt{X}\rangle\\ =\beta_{0}\langle\mathtt{X},\mathtt{X}\rangle-\langle\beta_{0}\mathcal{P}_{\mathfrak{a}}\mathtt{X},(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{B}^{-1}\mathcal{P}_{\mathfrak{a}}\mathtt{X}\rangle,\end{gathered}$
which is eq. 21. To show $d\mathfrak{q}$ is a submersion, let $\tilde{\mathtt{X}}\in\tilde{\mathfrak{m}}$ be the lift eq. 25 of $\mathtt{X}$, then
$\begin{gathered}\langle\tilde{\mathtt{X}},\tilde{\mathtt{X}}\rangle_{\tilde{\mathcal{G}},\vec{\beta}}=\beta_{0}\langle\mathtt{X}-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X},\mathtt{X}-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}\rangle\\ -\langle\beta_{0}\mathcal{B}(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})^{-1}(-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}),-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}\rangle\\ =\beta_{0}\langle\mathtt{X},\mathtt{X}\rangle-2\beta_{0}\langle\mathcal{P}_{\mathfrak{a}}\mathtt{X},(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}\rangle+\beta_{0}\langle(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X},(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}\rangle\\ -\langle\mathcal{B}\mathcal{P}_{\mathfrak{a}}\mathtt{X},(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}\rangle\\ =\beta_{0}\langle\mathtt{X},\mathtt{X}\rangle-\langle 2\beta_{0}\mathcal{P}_{\mathfrak{a}}\mathtt{X}-\beta_{0}(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}+\mathcal{B}\mathcal{P}_{\mathfrak{a}}\mathtt{X},(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}\rangle\\ =\beta_{0}\langle\mathtt{X},\mathtt{X}\rangle-\beta_{0}\langle\mathcal{P}_{\mathfrak{a}}\mathtt{X},(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}\rangle,\end{gathered}$
where between the third to last and the second to last expression, we factor out $\langle,(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{B}^{-1}\mathcal{P}_{\mathfrak{a}}\mathtt{X}\rangle$, then simplify to get to the last expression, which is the metric in eq. 18. We have previously showed $d\mathfrak{q}_{{}_{\operatorname{I}_{\tilde{\mathcal{G}}}}}$ maps $\tilde{\mathfrak{m}}$ bijectively onto $\mathfrak{g}$ with the inverse given by the lift eq. 25. In particular, $d\mathfrak{q}_{{}_{\operatorname{I}_{\tilde{\mathcal{G}}}}}$ is surjective, and the invariance of the metric shows it is surjective at any point in $\tilde{\mathcal{G}}$. Thus, $\mathfrak{q}$ is a submersion. Since eq. 21 corresponds to the nondegenerate form eq. 18 except for a singular pair $(\beta_{0},\mathcal{B})$, it corresponds everywhere by continuity. All the required identities hold, hence eq. 18 is naturally reductive. ∎
Theorem 2.
With the metric eq. 18, the Christoffel function of two tangent vectors $\xi,\eta$ at $X\in\mathcal{G}$ is given by
(27) $\begin{gathered}\Gamma(\xi,\eta)=-\frac{1}{2}(\xi X^{-1}\eta+\eta X^{-1}\xi)\\ +\frac{1}{2}X\left([(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(X^{-1}\xi),X^{-1}\eta]+[(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(X^{-1}\eta),X^{-1}\xi]\right).\end{gathered}$
The geodesic starting at $X\in\mathcal{G}$ with initial velocity $\xi\in T_{X}\mathcal{G}$ is given by
(28) $\gamma(t)=X\operatorname{expm}(t\left(-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(X^{-1}\xi)+X^{-1}\xi\right))\operatorname{expm}(t(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(X^{-1}\xi)).$
Set $\mathtt{V}=X^{-1}\xi$, let $\eta\in T_{X}\mathcal{M}$ be another tangent vector, then the parallel transport of $\eta$ along the geodesics $\gamma$ to $X_{t}=\gamma(t)$ with initial point and velocity $(X,\xi)$ is
(29) $\begin{gathered}\mathtt{T}(\gamma)_{0}^{t}\eta=X\operatorname{expm}(t\left(-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{V}+\mathtt{V}\right))\operatorname{expv}(t\mathrm{P}_{\mathtt{V}},X^{-1}\eta)\operatorname{expm}(t(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{V})\\ \text{ where for }\mathtt{B}\in\mathfrak{g},\mathrm{P}_{\mathtt{V}}:\mathtt{B}\mapsto\frac{1}{2}\left([\mathtt{B},\mathtt{V}]+[(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{V},\mathtt{B}]-[(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{B},\mathtt{V}]\right).\end{gathered}$
The map $\mathrm{P}_{\mathtt{V}}$ is antisymmetric in $\langle,\rangle_{\vec{\beta}}$, $\langle\mathrm{P}_{\mathtt{V}}\mathtt{B},\mathtt{C}\rangle_{\vec{\beta}}+\langle\mathrm{P}_{\mathtt{V}}\mathtt{C},\mathtt{B}\rangle_{\vec{\beta}}=0$ for $\mathtt{V},\mathtt{B},\mathtt{C}\in\mathfrak{g}$.
Proof.
We will assume $\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B}$ is invertible, then use continuity as before.
We start with the geodesics. For the geodesic at point $X\in\mathcal{G}$ with initial velocity $\xi$, we lift $X$ to $(X,\operatorname{I}_{\mathcal{A}})$. Then the geodesic $(X(t),A(t))$ on $\tilde{\mathcal{G}}$ using the horizontal lift of $X^{-1}\xi$ in eq. 25 is given by
$\left(X\operatorname{expm}\left(t(X^{-1}\xi-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(X^{-1}\xi))\right),\operatorname{expm}\left(t\left(-\beta_{0}^{-1}(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{P}_{\mathfrak{a}}(X^{-1}\xi)\right)\right)\right)$
which maps by the submersion $\mathfrak{q}$ to the geodesic eq. 28.
We will compute the connection at $\operatorname{I}_{\mathcal{G}}$ then use the invariance property. For the geodesic starting at $X(0)=\operatorname{I}_{\mathcal{G}}$ with initial velocity $\dot{X}(0)=\xi$ of the form $X(t)=\operatorname{expm}(t(-M_{1}+M_{2}))\operatorname{expm}(tM_{1})$ for $M_{1},M_{2}\in\mathfrak{g}$, we have $\dot{X}(t)=\operatorname{expm}(t(-M_{1}+M_{2})(M_{2})\operatorname{expm}(tM_{1})$ and
$\ddot{X}(0)=-\Gamma(\xi,\xi)=(-M_{1}+M_{2})M_{2}+M_{2}M_{1}=-[M_{1},M_{2}]+M_{2}^{2}.$
We thus get $\Gamma(\xi,\xi)=-\xi^{2}+[(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(\xi),\xi]$ and eq. 27 follows by polarization and left invariance.
For parallel transport, we also start with $X=\operatorname{I}_{\mathcal{G}}$ then use left invariance. From [39, Proposition 2.12], the lifted parallel transport from $(\operatorname{I}_{\mathcal{G}},\operatorname{I}_{\mathcal{A}})$ is $(\eta(t)^{\mathcal{G}},\eta(t)^{\mathcal{A}})=(\operatorname{expm}(t(\xi+\xi^{\mathcal{A}}))\mathtt{C}^{\mathcal{G}},\operatorname{expm}(t\xi^{\mathcal{A}})\mathtt{C}^{\mathcal{A}})$ where $(\mathtt{C}^{\mathcal{G}}(t),\mathtt{C}^{\mathcal{A}}(t))\in\mathfrak{g}\times\mathfrak{a}$ solves
$\begin{gathered}(\dot{\mathtt{C}}^{\mathcal{G}},\dot{\mathtt{C}}^{\mathcal{A}})=\mathcal{P}_{\tilde{\mathfrak{m}}}(-\frac{1}{2}[\xi+\xi^{\mathcal{A}},\mathtt{C}^{\mathcal{G}}],-\frac{1}{2}[\xi^{\mathcal{A}},\mathtt{C}^{\mathcal{A}}]),\\ (\mathtt{C}^{\mathcal{G}}(0),\mathtt{C}^{\mathcal{A}}(0))=(\eta+\eta^{\mathcal{A}},\eta^{\mathcal{A}}).\end{gathered}$
Since both the initial condition and the equation are in $\tilde{\mathfrak{m}}$, $(\mathtt{C}^{\mathcal{G}},\mathtt{C}^{\mathcal{A}})\in\tilde{\mathfrak{m}}$. Using eq. 24, the lifted transport maps to the transport on $\mathcal{G}$ as
$X(t)(X(t)^{-1}\eta(t)^{\mathcal{G}}-A(t)^{-1}\eta(t)^{\mathcal{A}})A(t)^{-1}=\operatorname{expm}(t(\xi+\xi^{\mathcal{A}}))(\mathtt{C}^{\mathcal{G}}-\mathtt{C}^{\mathcal{A}})\operatorname{expm}(-t\xi^{\mathcal{A}}).$
Let $F(t)=\mathtt{C}^{\mathcal{G}}(t)-\mathtt{C}^{\mathcal{A}}(t)$, thus, $(\mathtt{C}^{\mathcal{G}}(t),\mathtt{C}^{\mathcal{A}}(t))$ is the lift of $F(t)$ by $d\mathfrak{q}_{(\operatorname{I}_{\mathcal{G}},\operatorname{I}_{\mathcal{A}})}$. We have $F(0)=\eta$, and by eq. 25, $\operatorname{C}^{\mathcal{G}}=F+\operatorname{C}^{\mathcal{A}},\operatorname{C}^{\mathcal{A}}=-(\operatorname{I}_{\mathfrak{a}}+\beta^{-1}_{0}\mathcal{B})\mathcal{P}_{\mathfrak{a}}F$. The ODE for $F$ is
$\begin{gathered}\dot{F}=\dot{\mathtt{C}}^{\mathcal{G}}-\dot{\mathtt{C}}^{\mathcal{A}}=-\frac{1}{2}\left([\xi+\xi^{\mathcal{A}},F+\operatorname{C}^{\mathcal{A}}]-[\xi^{\mathcal{A}},\operatorname{C}^{\mathcal{A}}]\right)\\ =-\frac{1}{2}\left([\xi,F]-[\xi,(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}F]-[(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\xi,F]\right).\end{gathered}$
Thus, $F(t)=\operatorname{expv}(t\mathrm{P}_{\xi},\eta)$, giving us the parallel transport formula eq. 29 for $X=\operatorname{I}_{\mathcal{G}}$. To show $\mathrm{P}_{\mathtt{V}}$ (where $\mathtt{V}=\xi$) is antisymmetric in $\langle,\rangle_{\vec{\beta}}$, set $R=\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B}$, we calculate
$\begin{gathered}\frac{2}{\beta_{0}}\langle\mathrm{P}_{\mathtt{V}}\mathtt{B},\mathtt{C}\rangle_{\vec{\beta}}=\langle[\mathtt{B},\mathtt{V}]+[R\mathcal{P}_{\mathfrak{a}}\mathtt{V},\mathtt{B}]-[R\mathcal{P}_{\mathfrak{a}}\mathtt{B},\mathtt{V}],\mathtt{C}\rangle-\langle[\mathtt{B},\mathtt{V}]+[R\mathcal{P}_{\mathfrak{a}}\mathtt{V},\mathtt{B}]-[R\mathcal{P}_{\mathfrak{a}}\mathtt{B},\mathtt{V}],R\mathcal{P}_{\mathfrak{a}}\mathtt{C}\rangle\end{gathered}$
from eq. 23. Expand the right hand side and transform $\langle[R\mathcal{P}_{\mathfrak{a}}\mathtt{V},\mathtt{B}],R\mathcal{P}_{\mathfrak{a}}\mathtt{C}\rangle$ to
$\begin{gathered}\langle\mathtt{B},[R\mathcal{P}_{\mathfrak{a}}\mathtt{C},R\mathcal{P}_{\mathfrak{a}}\mathtt{V}]\rangle=\langle\mathcal{P}_{\mathfrak{a}}\mathtt{B},[R\mathcal{P}_{\mathfrak{a}}\mathtt{C},R\mathcal{P}_{\mathfrak{a}}\mathtt{V}]\rangle=\langle\mathcal{P}_{\mathfrak{a}}\mathtt{B},R[R\mathcal{P}_{\mathfrak{a}}\mathtt{C},\mathcal{P}_{\mathfrak{a}}\mathtt{V}]\rangle\\ =\langle R\mathcal{P}_{\mathfrak{a}}\mathtt{B},[R\mathcal{P}_{\mathfrak{a}}\mathtt{C},\mathcal{P}_{\mathfrak{a}}\mathtt{V}]\rangle=\langle\mathcal{P}_{\mathfrak{a}}\mathtt{V},[R\mathcal{P}_{\mathfrak{a}}\mathtt{B},R\mathcal{P}_{\mathfrak{a}}\mathtt{C}]\rangle=\langle\mathtt{V},[R\mathcal{P}_{\mathfrak{a}}\mathtt{B},R\mathcal{P}_{\mathfrak{a}}\mathtt{C}]\rangle.\end{gathered}$
Rearrange the other terms in $\frac{2}{\beta_{0}}\langle\mathrm{P}_{\mathtt{V}}\mathtt{B},\mathtt{C}\rangle_{\vec{\beta}}$ to isolate $\mathtt{V}$ to one side using bi-invariance,
$\begin{gathered}\frac{2}{\beta_{0}}\langle\mathrm{P}_{\mathtt{V}}\mathtt{B},\mathtt{C}\rangle_{\vec{\beta}}=\langle\mathtt{V},[\mathtt{C},\mathtt{B}]\rangle+\langle R\mathcal{P}_{\mathfrak{a}}\mathtt{V},[\mathtt{B},\mathtt{C}]\rangle-\langle\mathtt{V},[\mathtt{C},R\mathcal{P}_{\mathfrak{a}}\mathtt{B}]\rangle-\langle\mathtt{V},[R\mathcal{P}_{\mathfrak{a}}\mathtt{C},\mathtt{B}]\rangle\\ -2\langle\mathtt{V},[R\mathcal{P}_{\mathfrak{a}}\mathtt{B},R\mathcal{P}_{\mathfrak{a}}\mathtt{C}]\rangle\end{gathered}$
Thus, by anticommutativity of the Lie bracket, we have $\langle\mathrm{P}_{\mathtt{V}}\mathtt{B},\mathtt{C}\rangle_{\vec{\beta}}+\langle\mathrm{P}_{\mathtt{V}}\mathtt{C},\mathtt{B}\rangle_{\vec{\beta}}=0$. ∎
@@ 4.3. Riemannian metrics from the family
We will assume $\langle,\rangle$ is given by the trace form, and the algebras are transposable. The case $\mathcal{B}=-\beta_{0}\operatorname{I}_{\alpha}$ means the metric eq. 18 is bi-invariant, a multiple $\beta_{0}\operatorname{Tr}\mathtt{A}_{1}\mathtt{A}_{2}$ of the trace form. In theorem 2, the connection $\Gamma(\xi,\eta)$ reduces to the first two terms, the geodesic is $X\operatorname{expm}(tX^{-1}\xi)$, or just $\operatorname{expm}(t\xi)$ if $X=\operatorname{I}_{\mathcal{E}}$, the operator $\mathrm{P}_{\mathtt{V}}$ is just $\frac{1}{2}\operatorname{ad}_{\mathtt{V}}$, and using eq. 6, the parallel transport is $X\operatorname{expm}(\frac{t}{2}\mathtt{V})X^{-1}\eta\operatorname{expm}(\frac{t}{2}\mathtt{V})$. This is known [39].
However, the trace form metric may not be Riemannian. Using theorem 1 to deform the trace form using a Lie algebra $\mathfrak{a}$, for certain choices of the subgroup $\mathfrak{a}$ and parameters $\vec{\beta}$, the metric becomes Riemannian. Let $\mathsf{T}$ be the transpose operator, $\mathsf{T}:A\mapsto A^{\mathsf{T}}$. Compared with the Frobenius form, the deformation metric becomes
(30) $\langle\mathtt{A},\mathtt{B}\rangle_{\vec{\beta}}=\langle\mathtt{A},\mathcal{I}_{\vec{\beta}}\mathtt{B}\rangle_{\mathscr{F}},$
(31) $\mathcal{I}_{\vec{\beta}}:=\mathsf{T}\circ\left(\beta_{0}(\rm{I}_{\mathcal{E}}-\mathcal{P}_{\mathfrak{a}})-\mathcal{B}\circ\mathcal{P}_{\mathfrak{a}}\right)=\left(\beta_{0}(\rm{I}_{\mathcal{E}}-\mathcal{P}_{\mathfrak{a}})-\mathcal{B}\circ\mathcal{P}_{\mathfrak{a}}\right)\circ\mathsf{T}.$
In the above, we identify an element in the range of $\mathcal{B}$ with an element of $\mathfrak{g}$.
We will discuss a more general example in section 5, but will focus on the case $\mathcal{B}=\beta_{1}\operatorname{I}_{\mathfrak{a}}$ otherwise. The transpose $\mathsf{T}$ commutes with $\mathcal{P}_{\mathfrak{a}}$ since we consider transposable algebras. We will study the eigenvalues of $\mathcal{I}_{\vec{\beta}}=\mathsf{T}\circ\left(\beta_{0}(\rm{I}_{\mathcal{E}}-\mathcal{P}_{\mathfrak{a}})-\beta_{1}\mathcal{P}_{\mathfrak{a}}\right)$.
Note that the transpose $\mathsf{T}$ has eigenvalues $\pm 1$ on the spaces of symmetric and antisymmetric matrices. Write $\mathcal{V}_{\mathrm{sym}}$ and $\mathcal{V}_{\mathrm{skew}}$ for the symmetric and antisymmetric components of a transposable subspace $\mathcal{V}\subset\mathcal{E}$, the possible eigenpairs of $\mathcal{I}_{\vec{\beta}}$ are $(\beta_{0},(\mathfrak{a}_{{}_{\perp}})_{\mathrm{sym}}),(-\beta_{0},(\mathfrak{a}_{{}_{\perp}})_{\mathrm{skew}}),(-\beta_{1},\mathfrak{a}_{\mathrm{sym}}),(\beta_{1},(\mathfrak{a})_{\mathrm{skew}})$. The positive-definite requirement of a Riemannian metric implies at least one of the first two, and at least one of the last two eigenspaces are trivial, assuming the appropriate signs of $\beta_{0}$ and $\beta_{1}$. These are the possible cases
@@
    $(\mathfrak{a}_{{}_{\perp}})_{\mathrm{sym}}=\{0\},(\mathfrak{a})_{\mathrm{sym}}=\{0\}$. Thus, $\mathfrak{g}_{\mathrm{sym}}=\{0\}$, or $\mathcal{G}$ is a compact Lie subgroup of $\mathrm{SO}(n)$. In this case, $\beta_{0}<0,\beta_{1}>0$. An example is $\mathrm{SO}(n)$ itself in section 6.
@@
    $(\mathfrak{a}_{{}_{\perp}})_{\mathrm{sym}}=\{0\},(\mathfrak{a})_{\mathrm{skew}}=\{0\}$. For example, $\mathcal{G}$ could be a product of a compact subgroup (for $\mathfrak{a}_{{}_{\perp}}$) and a diagonal group (for $\mathfrak{a}$). We need $\beta_{0}<0,\beta_{1}<0$.
@@
    $(\mathfrak{a}_{{}_{\perp}})_{\mathrm{skew}}=\{0\},(\mathfrak{a})_{\mathrm{sym}}=\{0\}$. This shows $\mathfrak{a}=\mathfrak{g}_{\mathrm{skew}}$. An example is $\operatorname{GL}(n)$ in [42, 24], reviewed under our point of view with parallel transport in section 5. The case of $\mathrm{Sp}(n)$ in [8] could be treated similarly. In this case $\beta_{0}>0,\beta_{1}>0$. Except for the case $\mathfrak{g}_{\mathrm{sym}}=\{0\}$ already considered, the trace form is indefinite, it is positive on $\mathfrak{g}_{\mathrm{sym}}$ and negative on $\mathfrak{g}_{\mathrm{skew}}$. The deformation metric eq. 18 modifies the trace form on the antisymmetric part by a negative multiplier $\beta_{1}$. This is also known in [3].